OpenStudy (anonymous):

The volume of an inflated spherical water balloon is 288,000 mm3. What is the radius of the inflated water balloon? i get 41 ??

OpenStudy (anonymous):

nvm its 60 thanks

OpenStudy (anonymous):

It's not quite \(60\)...I think you forgot a factor of \(\pi\).\[V = \frac{4}{3}\pi r^3 \longrightarrow r = \sqrt[3]{\frac{3V}{4\pi}} \]I believe your first answer was correct (\(\approx 40.967\)).

OpenStudy (anonymous):

it is its cus sorry i forgt to write its really 288,000 pi mm^3 but u got it right if it werent for pi thaNKS ! =)